# Order-Robust Class Incremental Learning: Graph-Driven Dynamic Similarity Grouping

### v1

@@ Abstract
Class Incremental Learning (CIL) requires a model to continuously learn new classes without forgetting previously learned ones. While recent studies have significantly alleviated the problem of catastrophic forgetting (CF), more and more research reveals that the order in which classes appear have significant influences on CIL models. Specifically, prioritizing the learning of classes with lower similarity will enhance the model’s generalization performance and its ability to mitigate forgetting. Hence, it is imperative to develop an order-robust class incremental learning model that maintains stable performance even when faced with varying levels of class similarity in different orders. In response, we first provide additional theoretical analysis, which reveals that when the similarity among a group of classes is lower, the model demonstrates increased robustness to the class order. Then, we introduce a novel Graph-Driven Dynamic Similarity Grouping (GDDSG) method, which leverages a graph coloring algorithm for class-based similarity grouping. The proposed approach trains independent CIL models for each group of classes, ultimately combining these models to facilitate joint prediction. Experimental results demonstrate that our method effectively addresses the issue of class order sensitivity while achieving optimal performance in both model accuracy and anti-forgetting capability. Our code is available at https://github.com/AIGNLAI/GDDSG.
@@ 1 Introduction
Intra-task class conflicts refers to the discrepancies in model performance caused by similarity between classes that are trained simultaneously in a specific task (see Figure 1(b)). In real-world applications, where the arrival of classes in the data stream is uncontrollable, significant similarities among classes can severely impact the model’s resilience. For example, in a specific task from a sequence of tasks, a model may be trained to recognize different breeds within the same species. Within this task, due to the high similarity of features across categories, the model needs to develop resilience in distinguishing between closely related classes. However, existing CIL methods struggle to address this challenge, primarily due to the inherent limitations of the task setting. As CIL incrementally processes different classes, it cannot globally account for all class information, causing class conflicts to accumulate during training and negatively impact model performance. Thus, alleviating class conflicts and improving the model’s generalization ability remains a significant challenge in CIL.
@@ 3.2 The Effect of Class Ordering in CIL
Corollary 1 integrates the similarity between tasks with the model’s robustness to class order. Through Equation 1 and Equation 2, we observe that both forgetting and generalization errors are influenced by the optimal model gap between any two tasks, represented by $\lVert w^{*}_{i}-w^{*}_{j}\rVert^{2}$ for tasks $i$ and $j$. This gap serves as a measure of task similarity: the smaller the gap, the greater the similarity. Corollary 1 demonstrates that a smaller similarity between tasks enhances the model’s robustness in terms of generalization and resistance to forgetting across different class orders. This finding offers valuable insights for the design of new methods. The proof of Corollary 1 can be found in the supplementary material.
@@ 4 The Proposed Method: GDDSG
Figure 2: Illustration of The Overall Framework. [best view in color]
@@ 4.1 Class Grouping Based on Similarity
$\eta_{i,j}=\max[$ $\frac{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=i)\,d(x_{k}^{t},\mathbf{c_{i}})}{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=i)},$
$\frac{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=j)\,d(x_{k}^{t},\mathbf{c_{j}})}{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=j)}],$ (3)
where $d(\cdot,\cdot)$ denotes a distance metric function, $\mathbb{I}(\cdot)$ is an indicator function, and $\mathbf{c}_{i}=\frac{1}{|X|}\sum_{j=1}^{|X|}x_{j}$ represents the centroid of class $C_{i}$.
@@ 4.2 Graph-Driven Class Grouping
$P_{\text{Satisfy Brooks'}}=1-p^{2N}(1-p)^{N^{2}-2N}-p^{\frac{1}{2}N(N-1)}.$ (7)
Figure 3: A visualization of the region where the function $P_{\text{Satisfy Brooks}}$ is less than 0.99 is presented, with $p$ on the horizontal axis (ranging from 0.9 to 1.0) and $N$ on the vertical axis (ranging from 10 to 40). In regions not displayed, the corresponding $P_{\text{Satisfy Brooks}}$ values exceed 0.99.
@@
4.3 Overall Training Process
In the previous section, we introduced the motivation and core concepts behind the proposed algorithm. In this section, we will describe the entire training process in detail. Recent years have seen CIL methods based on pre-trained models achieve remarkable results [31, 59, 59, 27], largely due to their robust representation capabilities. Since our proposed class grouping method also relies heavily on the model’s representation ability, we utilize a widely-adopted pre-trained model as a feature extractor. For each class group, we train independent classification heads, which enhances the model’s adaptability and generalization to different class groups.
As outlined above, we utilize a frozen random projection matrix $W\in\mathbb{R}^{L\times M}$ to enhance features across all class groups, where $L$ is the output dimension of the pre-trained model and $M\gg L$ is the expanded dimensionality. Given a task $t$ and a sample $x^{t}_{i}$ belonging to a class group $s$, the feature vector of the sample is denoted as $h(x^{t}_{i})$, and its one-hot encoded label as $y(x^{t}_{i})$. Specifically,
$h(x^{t}_{i})=g(\phi(x)^{T}W),$ (8)
where $\phi(\cdot)$ represents the feature extractor, and $g(\cdot)$ is a nonlinear activation function. We define $H_{s}^{t}\in\mathbb{R}^{N_{s}^{t}\times M}$ as the matrix containing feature vectors of $N_{s}^{t}$ samples from group $s$. The corresponding Gram matrix is defined as:
$Gram_{s}^{t}={H_{s}^{t}}^{T}H_{s}^{t}\in\mathbb{R}^{M\times M}.$ (9)
Additionally, the matrix $C_{s}^{t}$ consists of the concatenated column vectors of all classes within group $s$, with dimensions $M\times L_{s}^{t}$, where $L_{s}^{t}$ represents the number of classes in group $s$ for task $t$. When a new task arrives, the model applies the GDDSG algorithm to assign new classes to their respective groups. The Gram matrix $Gram$ and matrix $C$ for each group are updated according to the following formulas:
$Gram_{s}^{t}=Gram_{s}^{t-1}+\sum_{n=1}^{N_{s}^{t}}h(x^{t}_{i})\otimes h(x^{t}_{i}),$ (10)
$C_{s}^{t}=\begin{bmatrix}C_{s}^{t-1}&\underbrace{\mathbf{0}_{M}\ \mathbf{0}_{M}\ \ldots\ \mathbf{0}_{M}}_{(L_{s}^{t}-L_{s}^{t-1})\text{ times}}\end{bmatrix}+\sum_{n=1}^{N_{s}^{t}}h(x^{t}_{i})\otimes y(x^{t}_{i}),$ (11)
where $\mathbf{0}_{M}$ denotes a zero vector with $M$ dimensions.
During the test phase, we combine the classification heads of all groups $G=[G_{1},G_{2},\dots,G_{k}]$ to make a joint prediction for a given sample $x$. For each class $c^{\prime}$ in a group, the score is computed as follows:
$s_{c^{\prime}}=g(\phi(x)^{T}W)(Gram_{i}+\lambda I)^{-1}C_{c^{\prime}},$ (12)
where $i=1,\dots,k$ denotes the indices of each groups, and $\lambda$ is the regularization parameter used to ensure that the $Gram$ matrix remains invertible. The final classification result is then obtained by applying the following formula:
$\hat{c}=\mathop{\arg\max}\limits_{c^{\prime}\in\cup_{i=1}^{k}CLS^{G_{i}}}s_{c^{\prime}},$ (13)
where $\cup_{i=1}^{k}CLS^{G_{i}}$ represents the set of possible classes across all class groups.
@@ 5 Experiment
Table 1: Results (%) of CL Methods on Both Fine-grained Datasets and General Vision Dataset. Among Them, The Best Results Are Bolded for Emphasis, While The Second-best Results Are Underlined.
@@
SimpleCIL 76.21 $\pm$ 0.00 7.51 $\pm$ 0.28 84.73 ±$\pm 0.45$ 5.01 $\pm$ 0.20 83.20 $\pm$ 0.88 5.85 $\pm$ 0.53 72.13 $\pm$ 0.00 8.49 $\pm$ 0.21
ADAM 83.66 $\pm$ 0.43 4.93 $\pm$ 0.08 85.35 $\pm$ 0.07 5.08 $\pm$ 0.29 84.21 $\pm$ 0.12 5.91 $\pm$ 0.59 72.63 $\pm$ 0.02 8.21 $\pm$ 0.19
EASE 87.19 $\pm$ 0.28 6.49 $\pm$ 0.31 88.76 $\pm$ 0.19 4.17 $\pm$ 0.27 80.06 $\pm$ 0.08 8.52 $\pm$ 0.14 77.62 $\pm$ 0.07 7.62 $\pm$ 0.28
RanPAC 90.74 $\pm$ 0.06 3.45 $\pm$ 0.19 89.48 $\pm$ 0.07 3.59 $\pm$ 0.04 85.10 $\pm$ 0.18 5.85 $\pm$ 0.50 78.77 $\pm$ 0.12 7.52 $\pm$ 0.12
GDDSG (Ours) 93.99 $\pm$ 0.03 0.72 $\pm$ 0.01 92.95 $\pm$ 0.00 1.01 $\pm$ 0.00 92.30 $\pm$ 0.01 1.48 $\pm$ 0.00 87.56 $\pm$ 0.01 1.07 $\pm$ 0.00
@@ 5.1 Experimental Setup
Implementations. Our code, implemented in PyTorch, has been open-sourced for accessibility. All experiments were conducted on a single Nvidia RTX 3090 GPU, using two random seeds, 2024 and 4202, to compute the average for a more robust model evaluation. We use a ViT-B/16 model, which is self-supervised and pre-trained on ImageNet-21K. Detailed dataset descriptions and experimental implementations are provided in the Supplementary Material.
@@
The Maximum OPD (MOPD) and Average OPD (AOPD) are further defined as:
$\text{MOPD}=\max\{\text{OPD}_{1},\ldots,\text{OPD}_{T}\},$ (15)
$\text{AOPD}=\frac{1}{T}\sum_{t=1}^{T}\text{OPD}_{t}.$ (16)
@@ 5.2 Experimental Results
Detailed Analysis of Class Group Matching. To better understand the mechanism of class group matching, we visualized all the distance features using t-SNE [38] and UMAP [28] on the Split CIFAR100 dataset. Figure 6 presents the visualized results. The observations reveal that the distance features essentially conform to piecewise functions in high dimensions, exhibiting strong linear separability and powerful representation capabilities. Consequently, a class group identification matching model can be effectively fitted using some classical machine learning models, enabling fairly accurate predictions. However, it is crucial to emphasize the importance of this step for GDDSG. Under the partition of the GDDSG algorithm, the accuracy of a single class group can reach nearly 100%. Therefore, the precision of class group matching directly determines the overall model’s accuracy.
@@ 6 Conclusion & Limitation
In this study, we aim to design an order-robust CIL model capable of addressing two critical challenges: class order sensitivity and intra-task conflicts. Building on existing theories, we find that as class similarity decreases, the model’s sensitivity to class order also lessens, which effectively mitigates knowledge conflicts both across tasks and within individual tasks. To enhance the model’s robustness across varying class orders, we propose a dynamic grouping method based on similarity graphs, termed GDDSG. The proposed approach maintains the centroids of learned classes and group classes based on dynamic similarity. In GDDSG, we introduce a novel approach to structuring class groups within class-incremental learning. Our GDDSG can continually update existing groups or form new ones, training distinct models for each group. During inference, predictions are derived through an ensemble of outputs from multiple models, thereby enhancing overall accuracy and robustness in CIL.
@@ Appendix A Notation
$SimG$ The graph generated base on 2

### v2

@@ Abstract
Class Incremental Learning (CIL) aims to enable models to learn new classes sequentially while retaining knowledge of previous ones. Although current methods have alleviated catastrophic forgetting (CF), recent studies highlight that the performance of CIL models is highly sensitive to the order of class arrival, particularly when sequentially introduced classes exhibit high inter-class similarity. To address this critical yet understudied challenge of class order sensitivity, we first extend existing CIL frameworks through theoretical analysis, proving that grouping classes with lower pairwise similarity during incremental phases significantly improves model robustness to order variations. Building on this insight, we propose Graph-Driven Dynamic Similarity Grouping (GDDSG), a novel method that employs graph coloring algorithms to dynamically partition classes into similarity-constrained groups. Each group trains an isolated CIL sub-model and constructs meta-features for class group identification. Experimental results demonstrate that our method effectively addresses the issue of class order sensitivity while achieving optimal performance in both model accuracy and anti-forgetting capability. Our code is available at https://github.com/AIGNLAI/GDDSG.
@@ 1 Introduction
Intra-task class conflicts refers to the discrepancies in model performance caused by the similarity between classes that are trained simultaneously in a specific task (see Figure 1(b)). In real-world applications, where the arrival of classes in the data stream is uncontrollable, significant similarities among classes can severely impact the model’s resilience. For example, in a specific task from a sequence of tasks, a model may be trained to recognize different breeds within the same species. Within this task, due to the high similarity of features across categories, the model needs to develop resilience in distinguishing between closely related classes. However, existing CIL methods struggle to address this challenge, primarily due to the inherent limitations of the task setting. As CIL incrementally processes different classes, it cannot globally account for all class information, causing class conflicts to accumulate during training and negatively impact model performance. Thus, alleviating class conflicts and improving the model’s generalization ability remains a significant challenge in CIL.
@@ 3.2 The Effect of Class Ordering in CIL
Corollary 1 integrates the similarity between tasks with the model’s robustness to class order. Through Equation 1 and Equation 2, we observe that both forgetting and generalization errors are influenced by the optimal model gap between any two tasks, represented by $\lVert w^{*}_{i}-w^{*}_{j}\rVert^{2}$ for tasks $i$ and $j$. This gap serves as a measure of task similarity: the smaller the gap, the greater the similarity. Corollary 1 demonstrates that a smaller similarity between tasks enhances the model’s robustness in terms of generalization and resistance to forgetting across different class orders. This finding offers valuable insights for the design of new methods.
@@ 4 The Proposed Method: GDDSG
Figure 2: Illustration of The Overall Framework. [best view in color]
@@ 4.1 Class Grouping Based on Similarity
$\eta_{i,j}=\max[$ $\frac{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=i)\,d(h(x_{k}^{t}),\mathbf{c_{i}})}{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=i)},$
$\frac{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=j)\,d(h(x_{k}^{t}),\mathbf{c_{j}})}{\sum_{k=1}^{|X^{t}|}\mathbb{I}(y^{t}_{k}=j)}],$ (3)
where $\mathbf{x}^{(t)}$ denotes the t-th task instance, $h(\cdot)$ is the feature extraction function defined in Equation Equation 8, $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{+}$ specifies the distance metric space, $\mathbb{I}(\cdot)$ represents the characteristic function, and the class centroid $\mathbf{c}_{i}\in\mathbb{R}^{d}$ is computed as $\mathbf{c}_{i}=\frac{1}{|C_{i}|}\sum_{x_{j}\in C_{i}}\mathbf{x}_{j}$.
@@ 4.2 Graph-Driven Class Grouping
$P_{\text{Satisfy Brooks}^{\prime}}=1-p^{2N}(1-p)^{N^{2}-2N}-p^{\frac{1}{2}N(N-1)}.$ (7)
Figure 3: Contour plot delineating the subthreshold region where $P_{\text{Satisfy Brooks}^{\prime}}<0.99$. The horizontal axis spans $p\in[0.9,1.0]$, representing probability values, while the vertical axis specifies sample sizes $N\in[10,40]$. In regions not displayed, the corresponding $P_{\text{Satisfy Brooks}^{\prime}}$ values exceed 0.99.
@@
4.3 Overall Process
Training Pipeline. Building upon the theoretical foundations in Section 3.1, we now formalize the complete training procedure. Our framework leverages a frozen pre-trained feature extractor $\phi(\cdot)$, augmented with trainable random projections $W\in\mathbb{R}^{L\times M}$ where $M\gg L$, to enhance representation capacity. For each input $x_{i}^{t}$ from class group $s$, we compute its expanded feature:
$h(x_{i}^{t})=g(\phi(x_{i}^{t})W)\in\mathbb{R}^{M},$ (8)
where $g(\cdot)$ denotes the nonlinear activation.
The core learning paradigm reframes classification as regularized least-squares regression. Let $H_{s}^{t}\in\mathbb{R}^{N_{s}^{t}\times M}$ be the feature matrix and $Y_{s}^{t}\in\mathbb{R}^{N_{s}^{t}\times L_{s}^{t}}$ denote the one-hot label matrix for class group $s$. We optimize the projection matrix $\Theta_{s}^{t}\in\mathbb{R}^{M\times L_{s}^{t}}$ through:
$\min_{\Theta}\|Y_{s}^{t}-H_{s}^{t}\Theta_{s}^{t}\|_{F}^{2}+\lambda\|\Theta_{s}^{t}\|_{F}^{2},$ (9)
where $\lambda$ controls regularization strength. The closed-form solution is:
$\Theta_{s}^{t}=({H_{s}^{t}}^{\top}H_{s}^{t}+\lambda I)^{-1}{H_{s}^{t}}^{\top}Y_{s}^{t}.$ (10)
For incremental updates, we maintain two key components: the Gram matrix $Gram_{s}^{t}$ capturing feature correlations, and the prototype matrix $C_{s}^{t}$ encoding class centroids. When new task $t$ arrives with $N_{s}^{t}$ samples:
$Gram_{s}^{t}=Gram_{s}^{t-1}+\sum_{n=1}^{N_{s}^{t}}h(x^{t}_{i})^{\top}h(x^{t}_{i}),$ (11)
$C_{s}^{t}=\begin{bmatrix}C_{s}^{t-1}&\underbrace{\mathbf{0}_{M}\ \mathbf{0}_{M}\ \ldots\ \mathbf{0}_{M}}_{(L_{s}^{t}-L_{s}^{t-1})\text{ times}}\end{bmatrix}+\sum_{n=1}^{N_{s}^{t}}h(x^{t}_{i})^{\top}y(x^{t}_{i}).$ (12)
The regularization parameter $\lambda$ is adaptively selected from a candidate pool $\Lambda$ through cross-validation on a held-out calibration set, minimizing the empirical risk:
$\lambda^{*}=\arg\min_{\lambda\in\Lambda}\|Y_{\text{val}}-H_{\text{val}}(Gram_{\text{val}}+\lambda I)^{-1}C_{\text{val}}\|_{F}^{2}.$ (13)
Additionally, group descriptors are constructed through prototype similarity analysis. For each training instance $(x,y)\in\mathcal{D}^{t}$, we generate meta-features dataset as:
$\mathcal{D}_{g}=\left\{\left(\rho(x),\ G^{-1}(y)\right)\right\}_{(x,y)\in\mathcal{D}^{t}},$ (14)
where $\rho(x)=\big{[}d(h(x),\mathbf{c}_{1}),\ldots,d(h(x),\mathbf{c}_{k})\big{]}^{\top}$ denotes the concatenated distance vector measuring similarity between the sample embedding and prototype vectors.
Inference Pipeline. Given test sample $x^{*}$, its group identification can be learned via $\hat{g}=\mathcal{M}_{g}(\rho(x^{*}))$, where $\mathcal{M}_{g}$ is the class group predict model trained with $\mathcal{D}_{g}$. Then, the prediction will be performed within the selected group via $\hat{y}=\underset{c\in\mathcal{C}_{\hat{g}}}{\arg\max}\ (g(\phi(x^{*})W)(Gram_{\hat{g}}+\lambda I)^{-1}C_{\hat{g}}[:,c]$.
@@ 5 Experiment
Table 1: Results (%) of CL Methods on Both Fine-grained Datasets and General Vision Dataset (Split into 10 Tasks). Among Them, The Best Results Are Bolded for Emphasis, While The Second-best Results Are Underlined.
@@
SimpleCIL 77.63 $\pm$ 2.01 7.21 $\pm$ 0.48 83.32 $\pm$ 2.03 5.89 $\pm$ 1.26 83.23 $\pm$ 0.72 5.86 $\pm$ 0.43 72.15 $\pm$ 0.02 8.31 $\pm$ 0.31
ADAM 84.15 $\pm$ 0.77 5.03 $\pm$ 0.15 83.89 $\pm$ 2.07 5.86 $\pm$ 1.12 83.98 $\pm$ 0.34 6.08 $\pm$ 0.54 72.84 $\pm$ 0.29 8.23 $\pm$ 0.16
EASE 86.95 $\pm$ 0.41 6.74 $\pm$ 0.43 85.78 $\pm$ 4.22 6.27 $\pm$ 2.98 81.03 $\pm$ 1.38 8.18 $\pm$ 0.50 74.56 $\pm$ 4.32 8.60 $\pm$ 1.40
RanPAC 90.50 $\pm$ 0.34 3.49 $\pm$ 0.16 89.23 $\pm$ 0.36 3.91 $\pm$ 0.45 85.37 $\pm$ 0.41 5.78 $\pm$ 0.42 78.86 $\pm$ 0.16 7.62 $\pm$ 0.18
GDDSG (Ours) 94.00 $\pm$ 0.03 0.78 $\pm$ 0.09 91.64 $\pm$ 1.86 1.92 $\pm$ 1.29 92.64 $\pm$ 0.48 1.42 $\pm$ 0.08 87.33 $\pm$ 0.32 0.96 $\pm$ 0.16
@@ 5.1 Experimental Setup
Implementations. Our code, implemented in PyTorch, has been open-sourced for accessibility. All experiments were conducted on a single Nvidia RTX 3090 GPU, using three random seeds, 1993, 2024, and 4202, to compute the average for a more robust model evaluation. We use a ViT-B/16 model, which is self-supervised and pre-trained on ImageNet-21K. $\mathcal{M}_{g}$ is a soft voting model consisting of RandomForest [4], KNN [35], and LightGBM [16].
@@
To characterize both extreme scenarios and overall performance, we introduce two consolidated metrics: Maximum OPD (MOPD) defined as $\text{MOPD}=\max\left\{\text{OPD}_{1},\ldots,\text{OPD}_{T}\right\}$, and Average OPD (AOPD) calculated by $\text{AOPD}=\frac{1}{T}\sum_{t=1}^{T}\text{OPD}_{t}$.
@@ 5.2 Experimental Results
Detailed Analysis of Class Group Identification. To better understand the mechanism of class group identification, we visualized $\mathcal{D}_{g}$ using t-SNE [41] and UMAP [30] on the Split CIFAR100 dataset. Figure 6 presents the visualized results. The observations reveal that the distance features essentially conform to piecewise functions in high dimensions, exhibiting strong linear separability and powerful representation capabilities. Consequently, a class group identification matching model can be effectively fitted using some classical machine learning models, enabling fairly accurate predictions. However, it is crucial to emphasize the importance of this step for GDDSG. Under the partition of the GDDSG algorithm, the accuracy of a single class group can reach nearly 100%. Therefore, the precision of class group matching directly determines the overall model’s accuracy.
@@ 6 Conclusion & Limitation
This study addresses the critical challenge of class order sensitivity in Class Incremental Learning (CIL), where model performance significantly degrades under varying class arrival sequences. By introducing GDDSG, a graph-driven framework that dynamically partitions classes into similarity-constrained groups and coordinates isolated sub-models with joint prediction, we theoretically and empirically mitigate the impact of class sequence variations. Experiments validate that our method not only reduces sensitivity to class order but also achieves state-of-the-art accuracy and anti-forgetting performance. This work provides a benchmark for developing robust CIL methods with dynamic data streams.
@@
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 62476228, No. 72242106), the Sichuan Science and Technology Program (No. 2024ZYD0180), and the Major Basic Research Project of Shandong Provincial Natural Science Foundation (No. ZR2024ZD03).
@@ Appendix A Notation
$SimG$ The graph generated base on Equation 4